OpenStudy (anonymous):

A projectile is fired straight up from ground level with an initial velocity of 112 ft/s. Its height, h, above the ground after t seconds is given by h = –16t^2 + 112t. What is the interval of time during which the projectile's height exceeds 192 feet? a. 3 < t < 4 b. t < 4 c. t > 4 d. 3 > t > 4

OpenStudy (anonymous):

So you set h to 192, then set it up as an inequality \[192ft <(112ft/s)t-(16ft/s^2)t^2\] So the 112t - 16t^2 has to be greater than 192 ft. Then you come up with the polynomial \[-(16ft/s^2)t^2+ (112ft/s)t-192ft > 0\] to make arithmetic simpler, that you can reduce to \[-(1ft/s^2)t^2+(7ft/s)t-(12ft)=0\] of the form \[at^2+bt+c>0\] and solve for t using the quadratic formula - time will be greater than the smaller root, and less than the larger root \[ t<\frac{-b-\sqrt{b^2-4ac}}{2a} \quad ; \quad t>\frac{-b+\sqrt{b^2-4ac}}{2a}\] \[a=-1 \\ b=+7 \\ c=-12\]

OpenStudy (anonymous):

so its D?